OpenStudy (anonymous):

What is the value of b in the equation below? 37 = 14b − 19 1 2/7 4 42 70 [Foil B] [Foil C] [Foil D]

OpenStudy (anonymous):

dont look at the foil thing

OpenStudy (amistre64):

avert thy gaze

OpenStudy (amistre64):

we need to get b all alone; so we need to get rid of the addons first

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

how do we make -19 go away?

OpenStudy (amistre64):

-19 + ? = 0

OpenStudy (anonymous):

just do it im writing down

OpenStudy (amistre64):

this is going to be short enough that your participation would be appreciated

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

you add both sides

OpenStudy (amistre64):

thats the right track yes; what should we add to get rid of the -19 ?

OpenStudy (anonymous):

37

OpenStudy (anonymous):

i think

OpenStudy (amistre64):

lets try it out -19 + 37 ?= 0 18 ?= 0 i dont think its 37

OpenStudy (anonymous):

14

OpenStudy (amistre64):

lets get rid of the -19 by adding 19 to both sides; keep in mind that we are trying to uncover the "b"

OpenStudy (amistre64):

37 = 14b − 19 +19 + 19 --------------- 56 = 14b + 0 56 = 14b

OpenStudy (anonymous):

oh yea i id that before

OpenStudy (amistre64):

now that the addons are taken care of; we need to get rid of the sticky part; that 14 is stuck to the "b" by multiplication to undo multiplication we divide

OpenStudy (amistre64):

computer froze :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yup we divide

OpenStudy (amistre64):

divide off the 14 to get to the "b" 56 = 14b /14 /14 --------- 56/14 = 1b and then we simplify 56/14 to match the an answer option

OpenStudy (anonymous):

4!

OpenStudy (anonymous):

THANK YOU

OpenStudy (amistre64):

youre welcome, and good luck :)